### v1

@@ 4. Conformal Prediction as a Weakly measurable Morphism
Weak measurability of $\kappa$. Let $\Lambda\coloneqq[0,1]\setminus S_{n+1}$. For every non-empty open set $G\subseteq Y$, put
$f_{G}(\alpha,y^{n},\psi)\;\coloneqq\;\sup_{y\in G}\pi(y,y^{n},\psi),\quad(\alpha,y^{n},\psi)\in\Lambda\times Y^{n}\times\mathscr{F}.$
Since $\pi(\cdot,y^{n},\psi)$ is upper semicontinuous and $G$ is open in the compact space $Y$, $f_{G}$ is itself upper semicontinuous and therefore Borel.
Because $\pi$ only takes values in $S_{n+1}$,
$\kappa(\alpha,y^{n},\psi)\cap G\neq\emptyset\quad\Longleftrightarrow\quad f_{G}(\alpha,y^{n},\psi)\;\geq\;\beta(\alpha).$
The map $\beta:\Lambda\to S_{n+1}$ is piecewise constant (hence Borel). Set
$\Theta_{G}\coloneqq\bigl{\{}(\alpha,y^{n},\psi)\in\Lambda\times Y^{n}\times\mathscr{F}:f_{G}(\alpha,y^{n},\psi)\geq\beta(\alpha)\bigr{\}}.$
Because $f_{G}$ is Borel and $\beta$ is Borel, $\Theta_{G}$ is a Borel subset of the product space. Therefore
$\bigl{\{}(\alpha,y^{n},\psi):\kappa(\alpha,y^{n},\psi)\cap G\neq\emptyset\bigr{\}}\;=\;\Theta_{G}\in\mathcal{B}\bigl{(}\Lambda\times Y^{n}\times\mathscr{F}\bigr{)}.$
In turn, for every open $G\subseteq Y$, the selector set $\{\kappa(\cdot)\cap G\neq\emptyset\}$ is measurable, hence the correspondence
$\kappa:\;\Lambda\times Y^{n}\times\mathscr{F}\rightrightarrows Y$
is weakly measurable (Aliprantis and Border,, 2006, Definition 18.1).
@@
    (i) Put total-variation on $\Delta_{Y}$. Then $P\mapsto P(A)$ is not continuous; Constraints in $\mathsf{CRED}$ are non-Borel, so weak measurability fails.
@@ 5. Unifying Bayesian, Frequentist, and Imprecise Reasoning
We call it $\mathsf{BCP}$ because it is evocative of an acronym for Bayesian Conformal Prediction (Fong and Holmes,, 2021). Next, we have an important lemma.
@@
We have $\lvert g\rvert\leq p(\theta)M(\theta)$ by (vii). Assumption (vi) gives continuity of $(y_{n+1},y^{n})\mapsto g(\theta,y_{n+1},y^{n})$ for $\lambda_{\Theta}$-almost all $\theta$. The envelope $pM^{n+1}\in L^{1}(\lambda_{\Theta})$ by (viii); Hence, dominated convergence yields continuity of $(y_{n+1},y^{n})\mapsto N_{P}(y_{n+1},y^{n})$.
@@
and similarly $\|D_{P_{m}}-D_{P}\|_{\infty}\to 0.$ Since $D_{P}$ is bounded below (by Step 2) and $D_{P_{m}}$ converges uniformly,
$\|p_{P_{m}}(\cdot\mid y^{n})-p_{P}(\cdot\mid y^{n})\|_{\infty}\xrightarrow[m\to\infty]{}0\quad\text{uniformly in }y^{n}.$
@@
As we showed before, $H$ is upper hemicontinuous in its three arguments, and $c(\alpha,y^{n},P)$ is continuous, hence he singleton-valued correspondence associated with $c$ (that we still denote by $c$ for ease of notation) is upper hemicontinuous. The composition of two upper hemicontinuous correspondences is again upper hemicontinuous (Aliprantis and Border,, 2006, Theorem 17.23). Therefore, $\mathsf{QUANT}$ is upper hemicontinuous. ∎
@@
First, notice that we do not require consonance directly because it is immediately satisfied as a result of (i) and $(y_{n+1},y^{n})\mapsto p_{P}(y_{n+1}\mid y^{n})$ being jointly continuous on $Y^{n}\times Y=Y^{n+1}$.
@@
Theorem 11 is the central payoff of our categorical treatment, and therefore the greatest joy (so far!) of looking at Conformal Prediction from a categorical angle. It confirms the intuition put forth in Martin, (2022) and Caprio et al., 2025a (, Section 5.2), that CP is a bridge between the Bayesian, frequentist, and imprecise approaches to predictive statistical reasoning. In the category UHCont, the model-based Bayesian and imprecise methods, and the model-free conformal construction, yield the same $\alpha$-level prediction region, thus showing a very profound connections between these three only apparently far apart prediction mechanisms.
@@
Suppose that $\overline{P}$ is $2$-alternating, that is, $\overline{P}(A_{1}\cup A_{2})+\overline{P}(A_{1}\cap A_{2})\leq\overline{P}(A_{1})+\overline{P}(A_{2})$, for all $A_{1},A_{2}\in\Sigma_{\Theta}$, and that $\sup_{y^{n}\in Y^{n}}\ell(y^{n}\mid\theta)<\infty$, for all $\theta\in\Theta$. Then, the upper posterior density
@@
where ¯$\underline{\ell}(y^{n})\coloneqq\int_{\Theta}\ell(y^{n}\mid\theta)\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)$, is an e-posterior (Grünwald,, 2023).
Proof.
First, notice that the upper posterior density in (10) is well defined if $\overline{P}$ is $2$-alternating and $\sup_{y^{n}\in Y^{n}}\ell(y^{n}\mid\theta)<\infty$, for all $\theta\in\Theta$, by Wasserman and Kadane, (1990); Caprio et al., 2024b .
@@
$\int_{\Theta}\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)\leq\overline{p}(\theta),\quad\forall\theta\in\Theta.$
But this corresponds to the coherence condition in Walley, (1991, Section 2.5), which is always met when $\mathcal{P}_{\text{prior}}\neq\emptyset$. The proof is then concluded. ∎
A consequence of Proposition 12 is that, if
$\overline{\Pi}(A)\equiv\overline{\Pi}_{(y^{n},\psi)}(A)=\int_{A}\underbrace{\int_{\Theta}\ell(y_{n+1}\mid\theta)\overline{p}(\theta\mid y^{n})\lambda_{\Theta}(\text{d}\theta)}_{\eqqcolon\overline{p}(y_{n+1}\mid y^{n})}\lambda_{Y}(\text{d}y),\quad\forall A\in\Sigma_{Y},$ (11)
@@
If (11) holds, then then there exist elements $y^{\ast}\in Y$ and $\theta^{\ast}\in\Theta$ such that
$\overline{p}(y_{n+1}\mid y^{n})=\delta_{y^{\ast}}(y)\quad(\lambda_{Y}\text{-a.s.}),\qquad\overline{p}\!\left(\theta\mid y^{n}\right)=\delta_{\theta^{\ast}}(\theta),\qquad\ell\!\left(y\mid\theta^{\ast}\right)=\delta_{y^{\ast}}(y).$
@@
Let $\overline{\Pi}_{\mathrm{B}}(A)\coloneqq\int_{A}\overline{p}(y_{n+1}\mid y^{n})\lambda_{Y}(\text{d}y)$, for all $A\in\Sigma_{Y}$.
Step 1. Maxitivity versus additivity. For disjoint measurable sets $A,B\in\Sigma_{Y}$,
$\overline{\Pi}(A\cup B)=\max\{\overline{\Pi}(A),\overline{\Pi}(B)\},\qquad\overline{\Pi}_{\mathrm{B}}(A\cup B)=\overline{\Pi}_{\mathrm{B}}(A)+\overline{\Pi}_{\mathrm{B}}(B),$
because $\overline{\Pi}$ is maxitive while $\overline{\Pi}_{\mathrm{B}}$ is additive. Equality (11) therefore forces
$\max\{\overline{\Pi}_{\mathrm{B}}(A),\overline{\Pi}_{\mathrm{B}}(B)\}\;=\;\overline{\Pi}_{\mathrm{B}}(A)+\overline{\Pi}_{\mathrm{B}}(B).$ (12)
Step 2. A 00-$1$ law for $\overline{\Pi}_{\mathrm{B}}$. If $10<\overline{\Pi}_{\mathrm{B}}(A)<$ 1 and $10<\overline{\Pi}_{\mathrm{B}}(B)<$ 1 with $A\cap B=\emptyset$, then (12) is impossible because $\max\{x,y\}<x+y$ for $x,y\in(0,1)$. Hence for every measurable $A$, it must be that
$\overline{\Pi}_{\mathrm{B}}(A)\in\{0,1\}.$ (13)
Step 3. Concentration on a single point. Define the set $S=\{y\in Y:\overline{p}(y_{n+1}\mid y^{n})>0\}$. If $\lambda_{Y}(S)=0$ then $\overline{\Pi}_{\mathrm{B}}\equiv 0$, contradicting $\overline{\Pi}_{\mathrm{B}}(Y)=1$. Thus $\lambda_{Y}(S)>0$ and (13) with $A=S$ gives $\overline{\Pi}_{\mathrm{B}}(S)=1$. Now take any two disjoint measurable subsets $A,B\subseteq S$ with positive $\lambda_{Y}$-measure. By (13) at most one of them can have $\overline{\Pi}_{\mathrm{B}}=1$, so we must have $\overline{\Pi}_{\mathrm{B}}(A)=\overline{\Pi}_{\mathrm{B}}(B)=0$. This is possible only if every measurable subset of $S$ with positive $\lambda_{Y}$-measure has $\overline{\Pi}_{\mathrm{B}}(\cdot)=1$, which forces $S$ to be a $\lambda_{Y}$-null neighborhood of a single point $y^{\ast}$ (up to $\lambda_{Y}$-null sets). Hence $\overline{\Pi}_{\mathrm{B}}$ (and therefore $\overline{p}(y_{n+1}\mid y^{n})$) is the point mass $\delta_{y^{\ast}}$, proving the first claim.
Step 4. Collapse of the posterior and the likelihood. Because
@@
the mixture representation can hold only if $\overline{p}(\theta\mid y^{n})=\delta_{\theta^{\ast}}(\theta)$ and $\ell(y\mid\theta^{\ast})=\delta_{y^{\ast}}(y)$. ∎
@@ 6. Conformal Prediction Regions as Images of the Imprecise Highest Density Region Functor
In this section, we see how the correspondence $\mathsf{IHDR}_{\alpha}$ can be seen as a functor between two well-known categories. As a consequence, the Conformal Prediction Region can be seen as a functor image.
@@
Let $(Y,\Sigma_{Y})$ be a measurable space, and let $\Delta_{Y}$ be the set of all finitely additive probability measures on $\Sigma_{Y}$ endowed with the weak topology. Let
@@ 7.1. Monoidal Categories
2. Associativity, unit, symmetry. Because $\otimes$ is the cartesian product on objects and the point-wise cartesian product on sets, we have equalities (not merely isomorphisms),
$(X\otimes Y)\otimes Z=X\otimes(Y\otimes Z),\qquad X\otimes\{\ast\}=X=\{\ast\}\otimes X.$
The associator and unitors are therefore identities, so the monoidal structure is strict. The symmetry is the flip map $\sigma_{X,Y}:(x,y)\mapsto(y,x)$; naturality and the hexagon condition hold because the cartesian product is symmetric in the category Set, and every construction is inherited verbatim.
3. Coherence. Since all coherence morphisms are identities, MacLane’s pentagon and triangle commute trivially.
Hence $(\mathbf{WMeas}_{\mathrm{uc}},\otimes,\{\ast\})$ is a strict symmetric monoidal category. ∎

### v2

@@ 4. Conformal Prediction as a Weakly measurable Morphism
Weak measurability of $\kappa$. Let $\Lambda\coloneqq[0,1]\setminus S_{n+1}$ and fix a non-empty open set $G\subseteq Y$. Because $Y$ is compact metric, it admits a countable dense set $Q\subseteq Y$; put $D_{G}\coloneqq Q\cap G$. Define
$f_{G}(\alpha,y^{n},\psi)\coloneqq\sup_{y\in D_{G}}\pi(y,y^{n},\psi),\qquad(\alpha,y^{n},\psi)\in\Lambda\times Y^{n}\times\mathscr{F}.$
Step 1: each coordinate map is Borel. Fix $y\in D_{G}$. For $i\in\{1,\dots,n+1\}$, let $g_{i}(y^{n},\psi)\coloneqq\psi(y^{n+1}_{-i},y_{i})-\psi(y^{n},y),$ which is continuous in $(y^{n},\psi)$ because evaluation of $\psi$ in the sup-norm is continuous and coordinate deletion is continuous. Hence, the set $\{(y^{n},\psi):g_{i}\geq 0\}$ is closed, its indicator is Borel, and
$(y^{n},\psi)\mapsto\pi(y,y^{n},\psi)=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbbm{1}[(y^{n},\psi):g_{i}\geq 0]$
is a finite sum of Borel functions, so Borel.
Step 2: the countable supremum is Borel. Because $D_{G}$ is countable, $f_{G}$ is the pointwise supremum of a countable family of Borel functions; Therefore $f_{G}$ itself is Borel.
Step 3: translate the selector set. Since $\pi$ takes values in the finite set $S_{n+1}$, define $\beta(\alpha)\coloneqq\min\{s\in S_{n+1}:s>\alpha\},$ a piecewise-constant (hence Borel) map $\beta:\Lambda\to S_{n+1}$. Then
$\kappa(\alpha,y^{n},\psi)\cap G\neq\emptyset\;\iff\;f_{G}(\alpha,y^{n},\psi)\geq\beta(\alpha).$
Consequently the selector set
$\Theta_{G}\coloneqq\bigl{\{}(\alpha,y^{n},\psi):f_{G}(\alpha,y^{n},\psi)\geq\beta(\alpha)\bigr{\}}$
is Borel in $\Lambda\times Y^{n}\times\mathscr{F}$. Thus
$(\kappa\cap G)^{-1}(\neq\emptyset)=\Theta_{G}\in\mathcal{B}\bigl{(}\Lambda\times Y^{n}\times\mathscr{F}\bigr{)},$
where $(\kappa\cap G)^{-1}(\neq\emptyset)\coloneqq\{(\alpha,y^{n},\psi):\kappa(\alpha,y^{n},\psi)\cap G\neq\emptyset\}$, proving that $\kappa$ is weakly measurable.
@@
    (i) Put Total Variation on $\Delta_{Y}$. Evaluation is still 1-Lipschitz $\bigl{(}|P(A)-Q(A)|\leq d_{\text{TV}}(P,Q)\bigr{)}$, but $\Delta_{Y}$ is not compact in TV. Hence the images $\mathsf{CRED}(y^{n},\psi)$ need not be compact (take the sequence of Dirac masses $\delta_{1/k}$ inside $\mathsf{CRED}$); Uniformly compact-valuedness fails.
@@ 5. Unifying Bayesian, Frequentist, and Imprecise Reasoning
We call it $\mathsf{BCP}$, an acronym for Bayesian Conformal Prediction (Fong and Holmes,, 2021). Next, we have an important lemma.
@@
We have $\lvert g\rvert\leq p(\theta)M(\theta)$ by (vii). Assumption (vi) gives continuity of $(y_{n+1},y^{n})\mapsto g(\theta,y_{n+1},y^{n})$ for $\lambda_{\Theta}$-almost all $\theta$. The envelope $pM\in L^{1}(\lambda_{\Theta})$ by (viii); Hence, dominated convergence yields continuity of $(y_{n+1},y^{n})\mapsto N_{P}(y_{n+1},y^{n})$.
@@
and similarly $\|D_{P_{m}}-D_{P}\|_{\infty}\to 0.$ Because $D_{P}$ is bounded below on the compact set $Y^{n}$ (by Step 2 there is a constant $\underline{d}_{P}\coloneqq\min_{y^{n}}D_{P}(y^{n})>0$) and $D_{P_{m}}$ converges uniformly, we have
$\inf_{y^{n}}D_{P_{m}}(y^{n})\geq\frac{1}{2}\,\underline{d}_{P}\quad\text{for all sufficiently large }m.$
Consequently
$\bigl{\|}p_{P_{m}}(\,\cdot\mid y^{n})-p_{P}(\,\cdot\mid y^{n})\bigr{\|}_{\infty}\xrightarrow[m\to\infty]{}0\quad\text{uniformly in }y^{n}\in Y^{n},$
and division remains well behaved thanks to the common lower bound $\tfrac{1}{2}\,\underline{d}_{P}$.
@@
As we showed before, $H$ is upper hemicontinuous in its three arguments, and $c(\alpha,y^{n},P)$ is continuous, hence the singleton-valued correspondence associated with $c$ (that we still denote by $c$ for ease of notation) is upper hemicontinuous. The composition of two upper hemicontinuous correspondences is again upper hemicontinuous (Aliprantis and Border,, 2006, Theorem 17.23). Therefore, $\mathsf{QUANT}$ is upper hemicontinuous. ∎
@@
First, notice that we do not require consonance directly because it is immediately satisfied as a result of (i) and $(y_{n+1},y^{n})\mapsto p_{P}(y_{n+1}\mid y^{n})$ being jointly continuous on $Y^{n}\times Y=Y^{n+1}$. They imply the existence of a maximizer $y^{\star}\in Y$ such that
$p_{P}(y^{\star}\mid y^{n})=\max_{y\in Y}p_{P}(y\mid y^{n}).$
Hence, $\psi_{(y^{n},P)}(y_{i})\geq\psi_{(y^{n},P)}(y^{\star})$, for every $i\in\{1,\dots,n\}$, and likewise for the candidate $y_{n+1}=y^{\star}$. Therefore every one of the $n\!+\!1$ indicator terms inside the transducer equals $1$, so
$\pi_{\psi_{(y^{n},P)}}(y^{\star},y^{n})=\frac{1}{n+1}\sum_{i=1}^{n+1}1=1.$
Consequently, $\sup_{y\in Y}\pi_{\psi_{(y^{n},P)}}(y,y^{n})=1$, that is, consonance holds automatically under (i) and the (joint) continuity of $p_{P}(y\mid y^{n})$.
@@
Theorem 11 is the central payoff of our categorical treatment, and perhaps the greatest joy (so far!) of looking at Conformal Prediction from a categorical angle. It confirms the intuition put forth in Martin, (2022) and Caprio et al., 2025a (, Section 5.2), that CP is a bridge between the Bayesian, frequentist, and imprecise approaches to predictive statistical reasoning. In the category UHCont, the model-based Bayesian and imprecise methods, and the model-free conformal construction, yield the same $\alpha$-level prediction region, thus showing a very profound connections between these three only apparently far apart prediction mechanisms.
@@
Suppose that $\underline{p}$ is integrable. Then, the upper posterior density
@@
where ¯$\underline{\ell}(y^{n})\coloneqq\int_{\Theta}\ell(y^{n}\mid\theta)\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)$, is an e-posterior à la Grünwald, (2023) if and only if $\int_{\Theta}\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)\leq\overline{p}(\theta)$, for all $\theta\in\Theta$.
We can interpret the necessary and sufficient condition in Proposition 12 as follows: The probability mass on which all priors agree is no larger than the minimum “peak plausibility” that any state enjoys under at least one prior.
Proof of Proposition 12.
First, notice that the upper posterior density in (10) is well defined. In particular, it is an upper bound for the generalized Bayes’ upper posterior density (Walley,, 1991, Theorem 6.4.6).
@@
$\int_{\Theta}\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)\leq\overline{p}(\theta),\quad\forall\theta\in\Theta.$ (11)
This concludes the proof. ∎
A consequence of Proposition 12 is that, if (11) holds and
$\overline{\Pi}(A)\equiv\overline{\Pi}_{(y^{n},\psi)}(A)=\int_{A}\underbrace{\int_{\Theta}\ell(y_{n+1}\mid\theta)\overline{p}(\theta\mid y^{n})\lambda_{\Theta}(\text{d}\theta)}_{\eqqcolon\overline{p}(y_{n+1}\mid y^{n})}\lambda_{Y}(\text{d}y_{n+1}),\quad\forall A\in\Sigma_{Y},$ (12)
@@
Assume (12). Then there exist points $y^{\ast}\in Y$ and $\theta^{\ast}\in\Theta$ such that, $\lambda_{Y}$- and $\lambda_{\Theta}$-almost surely,
$\overline{p}(y\mid y^{n})=\delta_{y^{\ast}}(y),\quad\overline{p}(\theta\mid y^{n})=\delta_{\theta^{\ast}}(\theta),\quad\ell(\,y\mid\theta^{\ast})=\delta_{y^{\ast}}(y).$
@@
Put
$\overline{\Pi}_{\mathrm{B}}(A)\coloneqq\int_{A}\overline{p}(y\mid y^{n})\,\lambda_{Y}(dy),\qquad A\in\Sigma_{Y},$
so (12) reads $\overline{\Pi}=\overline{\Pi}_{\mathrm{B}}$.
Step 1 (Maxitivity versus additivity). For disjoint $A,B\in\Sigma_{Y}$,
$\overline{\Pi}(A\cup B)=\max\{\overline{\Pi}(A),\overline{\Pi}(B)\},\qquad\overline{\Pi}_{\mathrm{B}}(A\cup B)=\overline{\Pi}_{\mathrm{B}}(A)+\overline{\Pi}_{\mathrm{B}}(B).$
Equality of $\overline{\Pi}$ and $\overline{\Pi}_{\mathrm{B}}$ forces
$\max\bigl{\{}\overline{\Pi}_{\mathrm{B}}(A),\overline{\Pi}_{\mathrm{B}}(B)\bigr{\}}\;=\;\overline{\Pi}_{\mathrm{B}}(A)+\overline{\Pi}_{\mathrm{B}}(B).$ (13)
Step 2 (A 00-$1$ law). If $\overline{\Pi}_{\mathrm{B}}(A),\overline{\Pi}_{\mathrm{B}}(B)\in(0,1)$, the right-hand side of (13) is strictly larger than the left, a contradiction. Hence
$\overline{\Pi}_{\mathrm{B}}(A)\in\{0,1\},\qquad\forall A\in\Sigma_{Y}.$ (14)
@@
[…]
implying $\overline{\Pi}_{\mathrm{B}}=\delta_{y^{\ast}}$.
Step 4 (Collapse of the posterior and the likelihood). The mixture representation
$\overline{p}(y\mid y^{n})=\int_{\Theta}\ell(y\mid\theta)\,\overline{p}(\theta\mid y^{n})\,\lambda_{\Theta}(d\theta)=\delta_{y^{\ast}}(y)$
can hold only if $\overline{p}(\theta\mid y^{n})=\delta_{\theta^{\ast}}(\theta)$ for some $\theta^{\ast}$ and $\ell(\,\cdot\mid\theta^{\ast})=\delta_{y^{\ast}}$. ∎
@@ 6. Conformal Prediction Regions as Images of the Imprecise Highest Density Region Functor
In this section, we see how the correspondence $\mathsf{IHDR}_{\alpha}$ can be cast as a functor between two well-known categories. As a consequence, the Conformal Prediction Region can be seen as a functor image.
@@
Let $(Y,\Sigma_{Y}=2^{Y})$ be a measurable space, and let $\Delta_{Y}$ be the set of all finitely additive probability measures on $\Sigma_{Y}$ endowed with the weak topology. Let
@@ 7.1. Monoidal Categories
2. Associativity, unit, symmetry. Because $\otimes$ is the cartesian product on objects and the point-wise cartesian product on fibres, we have the canonical homeomorphisms
$\alpha_{X,Y,Z}:(X\times Y)\times Z\;\longrightarrow\;X\times(Y\times Z),\qquad\lambda_{X}:\{\ast\}\times X\;\longrightarrow\;X,\qquad\rho_{X}:X\times\{\ast\}\;\longrightarrow\;X,$
and the symmetry (flip) map
$\sigma_{X,Y}:X\times Y\;\longrightarrow\;Y\times X,\;(x,y)\mapsto(y,x).$
Each of these maps is single-valued and continuous, hence a morphism in $\mathbf{WMeas}_{\text{uc}}$ (singleton fibres are compact). They play the role of associator, left- and right-unitors, and symmetry, respectively.
3. Coherence. The associator, unitors, and symmetry just introduced are the usual homeomorphisms that witness the cartesian product as a symmetric monoidal structure in Set. Consequently the pentagon, triangle, and hexagon coherence diagrams commute exactly as they do in Set, and the same proofs apply verbatim in $\mathbf{WMeas}_{\text{uc}}$.
Hence $(\mathbf{WMeas}_{\mathrm{uc}},\otimes,\{\ast\})$ is a symmetric monoidal category. ∎